OpenStudy (anonymous):

f(x)=x/(9+x^2), find power series. Any hints?

OpenStudy (blockcolder):

First find the power series (assuming it is centered at 0) of 1/(9+x^2). Then multiply each term by x.

OpenStudy (anonymous):

yes a swell idea.

OpenStudy (anonymous):

1/(1+x), now should i sub x=(x/3)^2,? i mean is it legal

OpenStudy (anonymous):

and for \[\frac{1}{9+x^2}\] start with \[\frac{\frac{1}{9}}{1+\frac{x^2}{9}}\]

OpenStudy (blockcolder):

@iHelp Yes, that's legal. :D

OpenStudy (anonymous):

sort of legal. idea is to make it look like \[\frac{1}{1+x}\]if you can

OpenStudy (anonymous):

that'll do it

OpenStudy (anonymous):

series n=1 to infin of (-1)^n((x/3)^2n)(x/9)? someone check for me plz ?

OpenStudy (anonymous):

yes i think so but you can clean it up because you have x twice

OpenStudy (anonymous):

you know you have \[\frac{1}{9}\sum(-1)^n(\frac{x}{3})^{2n} x\]

OpenStudy (anonymous):

so the other x should be included in the first one

OpenStudy (anonymous):

i c, thanks.

OpenStudy (anonymous):

\[\frac{1}{9}\sum(-1)^n\frac{x^{2n+1}}{3^{2n}}\] might look better although of course it is the same